OpenStudy (gabylovesyou):

Cylinder A has a radius of 10 inches and a height of 5 inches. Cylinder B has a volume of 750π. What is the percent change in volume between cylinders A and B? Cylinder B is 50% smaller than cylinder A. Cylinder B is 75% smaller than cylinder A. Cylinder B is 50% bigger than cylinder A. Cylinder B is 200% bigger than cylinder A. @waterineyes @hartnn i have no idea how to do this

OpenStudy (anonymous):

Can you find volume of Cylinder A?

OpenStudy (anonymous):

\(Volume_{cyl A} = \pi \cdot r^2 \cdot h\) Am I right?

OpenStudy (anonymous):

And keep \(\pi\) as it is..

OpenStudy (gabylovesyou):

yes..... volume is pi * r^2 * h pi * 10^2 * 5 = pi500

OpenStudy (anonymous):

Yep.. So, out of two, you know Cylinder B has greater volume? Right?

OpenStudy (gabylovesyou):

yes

OpenStudy (gabylovesyou):

so its between C and D

OpenStudy (anonymous):

\[\% Increase = \frac{Bigger - smaller}{smaller} \times 100 \quad \%\]

OpenStudy (anonymous):

Getting?

OpenStudy (gabylovesyou):

so 250 / 500 * 100%

OpenStudy (anonymous):

If you don't like this, then we have another method too..

OpenStudy (anonymous):

Yep..

OpenStudy (gabylovesyou):

.5 so 50% ?

OpenStudy (anonymous):

Simplify that..

OpenStudy (anonymous):

Yep..

OpenStudy (gabylovesyou):

thank you! :D

OpenStudy (anonymous):

You are welcome dear.. :)